A researcher claims that 13% of all motorcycle helmets have manufacturing flaws that could potentially cause injury to the wearer. A sample of 150 of these helmets revealed that 18 contained such defects.

1
Expert's answer
2022-06-07T11:21:18-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.13"

"H_a:p\not=0.13"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{\dfrac{18}{150}-0.13}{\sqrt{\dfrac{0.13(1-0.13)}{150}}}\approx-0.3642"

Since it is observed that "|z|=0.3642< 1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-0.3642)=0.715709," and since "p=0.715709>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.13," at the "\alpha = 0.05" significance level.